OpenStudy (anonymous):

In the lab, Trey has two solutions that contain alcohol and is mixing them with each other. He uses 3 times as much Solution A as Solution B. Solution A is 17% alcohol and Solution B is 12% alcohol. How many milliliters of Solution B does he use, if the resulting mixture has 378 milliliters of pure alcohol?

OpenStudy (anonymous):

lets call the amount of solution B \(x\) making the amount of solution A \(3x\) (since it is three times as much) then the amount of alcohol in solution A will be \(.17\times 3x\) and the amount in solution B will be \(.17x\)

OpenStudy (anonymous):

oops i mean the amount in solution B will be \(.12x\)

OpenStudy (anonymous):

since the total is \(378\) you can set \[.17\times 3x+.12\times x=378\] and solve for \(x\)

OpenStudy (anonymous):

might be easier to get rid of the annoying decimals and solve \[17\times 3x+12\times x=37800\]